a smelter ha 640 pounds of a 50% silver alloy and 200 pounds of a 20% alloy. how much of the 50% silver alloy must be mixed with all of the 20% alloy to give a 30% silver alloy?
Answer by stanbon(75887)

a smelter has 640 pounds of a 50% silver alloy and 200 pounds of a 20% alloy. how much of the 50% silver alloy must be mixed with all of the 20% alloy to give a 30% silver alloy?
---------
Equation:
silver + silver = silver
-------
0.50x + 0.20*200 = 0.30(x+200)
-------
50x + 20*200 = 30x + 30*200
20x = 10*200
x = 100 lbs (amt. of 50% alloy needed)
-------------------------
Cheers,
Stan H.
-------------------